OpenStudy (anonymous):

Integration please. I have to use algebraic substitution with this question but I really need help. I will type it below.

OpenStudy (anonymous):

\[integrate \sqrt4 / \sqrt x^3 + 4\]

OpenStudy (anonymous):

If I am correct I let u = x^3 +4 du/dx=3x^2 and dx=dx+du/3x^2

OpenStudy (anonymous):

\[\int\frac{\sqrt 4}{\sqrt{x^3}}\,dx\]?

OpenStudy (anonymous):

Is the 3 in the bottom inside or outside the \sqrt?

OpenStudy (callisto):

Is your question : ?